Determine the (OH), pH, and pOH of a solution with a fH+of 7.5 x 10-12 M at 25°C. on] = pH = pOH = Determine the (H+), pH, anDetermine the (H+), pH, and pOH of a solution with an [OH-] of 3.9 x 10-Mat 25 °C. [H+] = pH = pOH = Determine the [H+], [OH-Determine the H+], OH-], and pOH of a solution with a pH of 13.69 at 25 °C. [OH-] = 1 pOH = Determine the [HT], [OH-], and pHDetermine the (H], [OH-], and pH of a solution with a pOH of 9.04 at 25 °C. [H*] = [OH-] = pH = 0

I know this is a long question, but I promise that it is only one question in my homework.
Answer #1

[H*] = 7.5x10-12 M According to the ionic product of water, [OH-] = 1.0x10-14 7.5x10-12 = 1.3x10-3M pH = -log[H]=-log(1.3x10